OpenStudy (anonymous):

8th grade Math (easy) Medal & Fan .. systems of equations ..

OpenStudy (anonymous):

Katie and Pamela put some money into their money boxes every week. The amounts of money (y), in dollars, in their money boxes after certain amounts of time (x), in weeks, are shown by the equations below: Katie y = 30x + 20 Pamela y = 20x + 90 After how many weeks will Katie and Pamela have the same amount of money in their money boxes, and what will that amount be? (1 point) 6 weeks, $150 6 weeks, $230 7 weeks, $150 7 weeks, $230

OpenStudy (anonymous):

And please explain your answer please, you don't to take a whole long time... just tell me how you know thats the answer?

OpenStudy (anonymous):

@ganeshie8 @pooja195

OpenStudy (faiqraees):

Pamela's money = Katie's money 20x + 90 = 30x + 20 Solve the equation for x. Note: Since you weren't online that's why I couldn't explain you. But if you feel confused about any part let me know.

OpenStudy (anonymous):

I am & was online. Can you please help me with the answer? I still don't understand how to do it with what you gave me... @FaiqRaees

OpenStudy (faiqraees):

Okay so y is the amount of money and x is the number of weeks so y = 30x +20 means Money = 30(number of weeks) +20 Right?

OpenStudy (anonymous):

Yes and that means I would try the answers, like y= 30(6)+20 and also try y=30(7)+20 but what I got made nothing clear and wasn't any of my answers I could select.

OpenStudy (anonymous):

When I did y=30(6)+20 I got 200 and when I did y=30(7)+20 I got 230 @FaiqRaees

OpenStudy (faiqraees):

No you're going ahead. Lets just focuse what I am telling. So you understand what x and y means?

OpenStudy (anonymous):

Yes, x = amount of time and y= amount of money @FaiqRaees

OpenStudy (faiqraees):

Okay so they are asking when will pamela and katie will have the same amount of money. Which means they are asking when their y(money) values be same. Right?

OpenStudy (anonymous):

yes exactly @FaiqRaees

OpenStudy (faiqraees):

So can we say Y(money) of Pamela = Y(money) of Katie?